A golf ball is selected at random from a golf bag. If the golf bag contains 8 type A​ balls, 5 type B​ balls, and 3 type C​ balls, find the probability that the golf ball is not a type A ball.


1
Expert's answer
2022-03-31T05:58:43-0400

"P(not A)=\frac{5+3}{8+5+3}=\frac{8}{16}=\frac{1}{2}."